OpenStudy (anonymous):

3c+24-2c=-9 (Show Work)

OpenStudy (bahrom7893):

3c-2c+24=-9 c = -9 -24 c = -33

OpenStudy (anonymous):

3c-2c+24=-9 c = -9 -24 c = -33

OpenStudy (anonymous):

You want to have the c on one side and numbers on another side of 3c+24-2c=-9 To do this, when you move a number to the other side of an equal sign, then you change the sign. A negative will be positive and a positive will be negative. So it would be 3c-2c=-24-9. which would be c=-33

OpenStudy (anonymous):

The 3c-2c is like this because when you move the 24, the + is part of the 24, but when it goes to the other side then it becomes negative. The -2 sign keeps it a negative.